OpenStudy (jessie13):

What is the surface area of a rectangular prism that has a length of 18 inches, a width of 1 foot, and a height of 14 inches? 252 in. 2 3,024 in. 2 1,272 in. 2 568 in. 2

OpenStudy (jessie13):

need help asap!

OpenStudy (automaticloveletter):

SA=2(wl+hl+hw) L=18 W=1 H=14 Plug in the numbers SA=2*(1*18+14*18+14*1) SA =568

OpenStudy (jessie13):

okay thankz